OpenStudy (anonymous):

I got answers idk which is correct though. In Oregon, 3/5 of bottles are recycled each year. A bottling company produces 1.2 million bottles and figures that x years later the number U of those sill being used is given by U(x)=1200270(3/5)^x Of those 1.2 million bottles, how many are still in use 8 years later? A. 20160 B. 9873 C. 1126024 and D. None of the

OpenStudy (anonymous):

i am not sure is the andser is 20160 or 2520. I have set U(x)=1200270(3/5)^x to look like U(8)=1200270(3/5)^8 and calculated for 2520. However, i am not sure if you have to divide by 8 on both sides because if I didnt i would get 20160

OpenStudy (anonymous):

i got U(8) =20159.9 so the answer is A

OpenStudy (anonymous):

thanks